Shelia's measured glucose level one hour after a sugary drink varies according to the normal distribution with ?=120 mg/dl and ?=10.7 mg/dl. What is the level ? such that there is probability only 0.03 that the mean glucose level of 2 test results falls above ?? Give your answer precise to one decimal place. ?=
Answer #1
mean μ= 120
standard deviation σ= 10.7000
sample size       =n= 2
std error=σ=σ/√n= 7.566
for 97th percentile critical value of z= 1.88
therefore corresponding value=mean+z*std deviation= 134.2